# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v2

@@ 5.2 RQ2: In-world visits of poker casinos
When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figurge 7) and number of distinct visitor accounts (see Figure 8). For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors, with a trend strength (based on STL decomposition) of 0.98 and 0.99, respectively. In the rest of the world map, the trend for time spent is slightly downward and with a strength of 0.83 for time spent, while the trend is visibly downward with a strength 0.91 for visitors. In other words, over time there were less distinct accounts visiting Decentraland, but these collectively spent more time on it. However, upon further inspection this increase in time spent per visitor is mostly due to a few accounts with a much higher number of hours than average. Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due the the aforementioned outlier users in the rest of the world map. Nevertheless, for visitors, the share at the beginning of the ten-month period was of circa 32%, and at the end it increased to circa 41%. Curiously, the number of visitors of ICE Poker casinos had a modest weekly seasonality strength of 0.66, with the number of visitors slightly decreasing on weekends; for the other time series the seasonality strength was less than 0.5.
@@
Figure 7: Stacked area chart of daily time spent (in hours) visiting Decentraland. ICE Poker casinos had a remarkable share, with a daily median of 20%. The highest overall spike was reached during the Metaverse Music Festival (MVMF) in mid-November.
@@
Figure 8: Stacked area chart of daily distinct accounts that visited Decentraland. ICE Poker casinos had a striking share, with a daily median of 33%. The Metaverse Music Festival (MVMF) in mid-November had the highest spike overall, while the Metaverse Fashion Week (MVFW) at the end of March had the only spike during the last four months.
@@
These trend and seasonality strength differences and shares by land category are in part explained by their dissimilar spikes and surges. On one hand, visits to the ICE Poker casinos had no spikes nor dips. There was, however, a short-lived surge (for both time and accounts) that started at the end of August 2022, when DG gifted special wearables to circa 3,000 of the most active players as a reward and in preparation for its new Tournament game mode. On the other hand, in the rest of the world map there were several visit spikes, surges, and dips. In particular, the Decentraland Metaverse Music Festival, held on 10–13 November 2022, was the event that provoked the most visits (both in time and accounts). The festival was free to attend, and its line up involved famous artists such as Björk, Ozzy Osbourne, Dillon Francis, and Soulja Boy. Incidentally, the Metaverse Fashion Week of 2023 also provoked a spike in visitors at the end of March, but not in time spent and only during a period of four months in which the number of distinct visitors had already remarkably declined.
6 Discussion
@@
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we first discuss our main findings concerning the influence of DG and ICE Poker on Decentraland, we comment on the initiatives and events affecting such influence over time, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and, finally, we describe the main limitations of our study.
6.1 Main findings
@@
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted (i.e., involved in a creation transfer) based on the maximum supply limit compared to the rest, particularly at the end of the period of study. DG wearables also had a monetary sales volume an order of magnitude higher despite the considerably fewer sales. Moreover, in spite of covering less that 0.1% of the land, ICE Poker casinos had a striking share of time spent and unique visitors in Decentraland, with a respective daily median of 20% and 33%. Further, the vast majority of these visits were done to a single casino: The Stronghold. Still, these dynamics have varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
@@
The Metaverse Fashion Week is arguably the most important initiative around Decentraland wearables. Both editions of 2022 and 2023 attracted well-established fashion brands, were featured in mainstream media, and piqued the interest of the general public. The initiative, however, did not meet expectations. Based on a qualitative study on the 2023 edition Miikkulainen-Gilbert (2023), most user complaints were related to the on-boarding complexity of setting up an Ethereum account, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction for many events, the technical limitations of wearable 3D modeling, and the need of paid approval even for wearables designed for oneself. In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure 8), but there was no visible effect on the time spent on the platform (see Figure 7), and the number of visitors continued its downward trend afterward. The fact that in 2024 there was no Metaverse Fashion Week confirms the failure of the initiative to keep up the momentum going. On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures 7 and 8), but nevertheless visits did not visibly increase in the following weeks. Unlike in-world fashion events, music concerts have been present for many years in social virtual worlds and MMO games, and during the Covid-19 pandemic they were seen as an attractive alternative for the struggling live music industry Çeli̇k (2023). It remains to be seen if this Decentraland initiative will continue to gather interest in future editions.
@@
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation. Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023). In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however. This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
@@
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023). Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6. Nevertheless, the initial hype regarding NFTs has subsided (and to a lesser extent also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT Echauri (2024).
In addition, platforms based on tokenomics must continually adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations OECD (2022). Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024). DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
6.2 On the legality of ICE Poker
@@
After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions Zborowska et al. (2012). Similarly, for many years and due to their novelty, blockchain cryptocurrencies remained in a regulatory limbo in most jurisdictions Yeoh (2017). However, after many instances of scams and fraud —some very large and highly mediatized, as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies Trautman et al. (2022). ICE Poker is a combination of both gambling and cryptocurrencies, hence, this begs the simple question on whether it is a legal operation or not. The answer is, however, murky.
@@
Firstly, on the project’s front page it is stated that DG is operated by the company Kidderminster Cheltenham, whose official address is in Costa Rica. This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  Kerr (2018), its tax haven tendencies Baron (2019), and the fact that its operating licenses are often used to offer an online gambling service in foreign jurisdictions in which the licensing process is more rigid or non-existing Fischer (2021). Moreover, in the official project documentation it is stated that the company behind DG is called Web4 LTD, which is established in the British Virgin Islands, the place of jurisdiction regarding the terms and conditions of the project, and a British overseas territory known for being a tax haven and having links to dubious online gambling operations Macegoniuk (2021).
Curiously, in the project’s website there is also a legal opinion regarding ICE Poker by a consultant from the Singaporean firm Jacque Law Zhirong (2022), in which it is stated that the game “would not constitute ‘gambling’, ‘betting’, ‘lottery’ under the RGA [Singaporean Remote Gambling Act of 2014]” because “gambling laws are not intended to cover games whereby there is no staking or wagering of money or money’s worth”. Indeed, possession of a DG wearable is necessary to earn ICE, but the wearable has independent value outside of ICE Poker and its use is not exclusive to DG. In addition, users play with chips, which according to the opinion Zhirong (2022), “are non-transferable tokens that hold no value or functionality other than to keep score when players play ICE Poker”. Based on the number of chips and wearables they possess, players periodically receive ICE, which can be converted to fiat money, but on an external exchange such as Binance. This convoluted mechanism is similar to the one used offline by pachinko parlours in Japan to circumvent gambling laws Brooks et al. (2008): pachinko players win small steel balls, which can be exchanged for a receipt with a nominal value, which in turn is used to claim a low-valued physical item (e.g., a specific pack of washing powder), and finally, the player can exchange this item for a lesser nominal value in cash outside the parlour on a nearby “independent office”.
Of course, the above legal opinion applies only to Singapore. In the case of other jurisdictions DG is more cautious, claiming that the onus is on players. In fact, to participate in its virtual casinos, players must guarantee they are not residents of one of the non-supported regions listed on the project’s website. At the moment of writing (early 2024), there were 36 non-supported regions, including high-income countries such as the United States, the United Kingdom, France, Germany, and Australia. ICE Poker Arcade might nonetheless be supported in these regions, given that users play to win physical or digital prizes, not money or cryptocurrency. To address the issue in a more solid manner, recently DG acquired a 90% stake in Virtue Gaming Ltd, the entity that holds a Malta license to enable real-money gambling in Virtue Poker, the first licensed cryptocurrency-based online poker. DG intends to expand this coveted license to its virtual casinos and other games, which would remove doubts regarding their current legal status in many jurisdictions.
6.3 Social implications
@@
Unlike most previous social virtual worlds and social media in general, platform transparency and openness are stated to be paramount in Decentraland. The project strives toward these qualities by means of its use of public records on the blockchain, user-oriented ownership of in-world assets, public and open-source code repositories, governance via a DAO, and initiatives to overtly monitor the ecosystem —e.g., by financing the collection of in-world tracking data we used to analyze visits. These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems, in the form of virtual reality spaces but not only Cannavo & Lamberti (2020). Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform’s mechanisms.
@@
For instance, in June 2022 DG submitted a proposal to Decentraland’s DAO for a grant, [^5] https://decentraland.org/governance/proposal/?id=99c66b90-e2d2-11ec-9000-175d8dd584b8 asking the equivalent to one million USD in MANA to increase its liquidity pool for ICE rewards. The grant was approved by 56% of the voting power, which is based on ownership of Decentraland MANA, usernames, and land holdings. DG —one of the earliest and most prominent actors in the platform— accounted for almost a third of the voting power in its own favor. This is an illustrative example of how a decentralized governance is not necessarily a neutral one. Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem Goldberg & Schär (2023), because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
@@
Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the current revamping of ICE Poker itself. These highlight that despite the assurances of public transactions and smart contracts, most of the functionality of the poker system —and the virtual world itself— lives outside the blockchain, and is thus subject to arbitrary changes by its controlling entity. In addition, a third-party actor was able to exploit the affordances provided by the virtual world, which lead to deviations in user behavior unforeseen by the platform’s creators. The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token). Hence, functional wearables might be appreciated solely for their utility, whereas cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a creative social space.
Besides, based on our results in-world visits seem be greatly influenced by the earning potential within ICE Poker casinos, instead of the ludic enjoyment that these and other world areas offer. However, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games Delic & Delfabbro (2024); Hong et al. (2023). At the same time, in online poker the earning prospects are partially integrated into the gameplay, thus becoming part of the intrinsic motivation to play, but with the possibility of leading to problem gambling without the proper containment mechanisms Zaman et al. (2014).
ICE Poker sits at the intersection of gamblers and cryptoasset traders, with potentially troubling consequences as both profiles might exhibit similar self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction Johnson et al. (2023); Delfabbro et al. (2021b). We should not forget, however, that in Ice Poker, as with other play-to-earn games, a considerable amount of grunt play is done by low-income people who do not own assets and for whom game enjoyment is secondary. It is a new form of digital exploitation —where the lion’s share goes to the owner of the NFT who lends it to the less fortunate— and which could lead to increased social disparity and potentially exacerbate economic inequalities Zaucha (2024).
For all the above reasons, the study of the emerging phenomena in novel social platforms such as Decentraland and games such as Ice Poker is increasingly important. A better understanding of their underlying mechanisms and dynamics will allow, for instance, polycymakers to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years.
6.4 Limitations

### v3

@@ B.   RQ2: In-world visits of poker casinos
When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figure VII) and number of distinct visitor accounts (see Figure VIII). For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors. In the rest of the world map, the trend is slightly downward for time spent, while the trend is visibly downward for visitors. In other words, over time there were less distinct accounts visiting Decentraland, but these collectively spent more time on it. However, upon further inspection this increase in time spent per visitor is mostly due to a few accounts with a much higher number of hours than average. Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due to the aforementioned outlier users in the rest of the world map. Nevertheless, for visitors, the share at the beginning of the ten-month period was of circa 32%, and at the end it increased to circa 41%.
@@
Figure VII: Daily time spent (in hours) visiting Decentraland. ICE Poker casinos had a remarkable share, with a daily median of 20%. The highest overall spike was reached during the Metaverse Music Festival (MVMF) in mid-November.
@@
Figure VIII: Daily distinct accounts that visited Decentraland. ICE Poker casinos had a striking share, with a daily median of 33%. The Metaverse Music Festival (MVMF) in mid-November had the highest spike overall, while the Metaverse Fashion Week (MVFW) at the end of March had the only spike during the last four months.
@@
On one hand, visits to the ICE Poker casinos had no spikes nor dips. There was, however, a short-lived surge (for both time and accounts) that started at the end of August 2022, when DG gifted special wearables to circa 3,000 of the most active players as a reward and in preparation for its new Tournament game mode. On the other hand, in the rest of the world map there were several visit spikes, surges, and dips. In particular, the Decentraland Metaverse Music Festival, held on 10–13 November 2022, was the event that provoked the most visits (both in time and accounts). Incidentally, the Metaverse Fashion Week of 2023 also provoked a spike in visitors at the end of March, but not in time spent and only during a period of four months in which the number of distinct visitors had already remarkably declined.
VI.   Discussion
@@
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we discuss our main findings, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and finally, we describe the main limitations of our study.
A.   Main findings
@@
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted based on the maximum supply limit compared to the rest, particularly at the end of the period of study. DG wearables also had a monetary sales volume an order of magnitude higher despite the considerably fewer sales. Moreover, in spite of covering less that 0.1% of the land, ICE Poker casinos had a striking share of time spent and unique visitors in Decentraland, with a respective daily median of 20% and 33%. Further, the vast majority of these visits were done to a single casino: The Stronghold. Still, these dynamics have varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
@@
The Metaverse Fashion Week is arguably the most important initiative around Decentraland wearables. Both editions of 2022 and 2023 attracted well-established fashion brands, were featured in mainstream media, and piqued the interest of the general public. The initiative, however, did not meet expectations. Based on a qualitative study on the 2023 edition [38], most user complaints were related to on-boarding complexity, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction, and the need of paid approval even for wearables designed for oneself. In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure VIII), but there was no visible effect on the time spent on the platform (see Figure VII), and the number of visitors continued its downward trend afterward. The fact that in 2024 there was no Metaverse Fashion Week confirms the failure of the initiative to keep up the momentum going. On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures VII and VIII), but nevertheless visits did not visibly increase in the following weeks. It remains to be seen if this initiative will continue to gather interest.
@@
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, accelerating its inflation. In order to not alienate core delegates with these changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures VII and VIII). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow players to earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however. This may explain the increase in the share of transfers of DG wearables at the end of the period of study (see Figure V) without the corresponding increase in visits share to ICE Poker casinos.
@@
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof [13]. Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures IV and VI. Nevertheless, the initial hype regarding NFTs has since subsided (also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT [15].
B.   On the legality of ICE Poker
@@
After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions [57]. Similarly, for many years and due to their novelty, cryptoassets remained in a regulatory limbo in most jurisdictions, but after many instances of scams and fraud —as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies [50]. ICE Poker is a combination of both gambling and cryptocurrencies. Hence, this begs the simple question on whether this kind of operation is legal or not. The answer is, however, murky.
@@
Firstly, on the project’s front page it was stated that DG’s parent company operated in Costa Rica. This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  [34], and the fact that its operating licenses are often used for online gambling services in foreign jurisdictions in which licensing is more rigid or non-existing [19]. Moreover, in the official project documentation it is stated that the company developing DG, Web4 LTD, is established in the British Virgin Islands, a British overseas territory known for being a tax haven and having links to dubious online gambling operations [37].
Curiously, the terms and conditions of ICE Poker state that is not covered by gambling laws in many jurisdictions due to the lack of (fiat) money exchange therein. Based on the number of chips and wearables they possess, players periodically receive ICE, which can ultimately be converted to fiat money, but on an external exchange such as Binance. Yet, players must guarantee they are not residents of one of the dozens of non-supported regions listed on the project’s website due to their strict gambling and cryptocurrency regulations.
C.   Social implications
@@
Platform transparency and openness are stated to be paramount in Decentraland. The project strives toward these qualities by means of its use of public records on the blockchain, user-oriented ownership of in-world assets, public and open-source code repositories, governance via a DAO, and initiatives to overtly monitor the ecosystem —e.g., by financing the collection of in-world tracking data we used to analyze visits. These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems in the form of virtual reality spaces, but not only [9]. Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform.
@@
For instance, in June 2022 DG submitted a proposal to Decentraland’s DAO for a grant, [^5] https://decentraland.org/governance/proposal/?id=99c66b90-e2d2-11ec-9000-175d8dd584b8 asking the equivalent to one million USD in MANA to increase its liquidity pool for ICE rewards. The grant was approved by 56% of the voting power, which is based on ownership of Decentraland MANA, usernames, and land holdings. DG —one of the earliest and most prominent actors in the platform— accounted for almost a third of the voting power in its own favor. This is an illustrative example of how a decentralized governance is not necessarily a neutral one. Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem [22], because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the revamping of ICE Poker itself. These highlight that despite the assurances of public transactions and smart contracts, most of the functionality of the poker system —and the virtual world itself— lives outside the blockchain, and is thus subject to arbitrary changes by its controlling entity.
@@
In addition, a third-party actor was able to exploit the affordances provided by the virtual world, which lead to deviations in user behavior unforeseen by the platform’s creators. The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, perhaps tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token). Hence, functional wearables might be appreciated solely for their utility, whereas merely cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a social space primarily focused on creativity and self-expression. In fact, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games [29].
Based on our results, in-world visits are also greatly related to DG casinos. As with wearable transfers, this phenomenon is likely influenced by the earning potential of ICE Poker, in addition to the ludic enjoyment alone that casinos and other world areas offer. Moreover, ICE Poker combines gambling and cryptoasset trading, with potentially troubling consequences to players due to self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction [31].
Lastly, professional gamblers in ICE Poker and ICE Arcade could potentially replicate exploitative practices observed in other Play-to-Earn (P2E) games. They might use scholarships and delegation as a form of digital feudalism, where the majority of the profits are funneled to a small group of individuals or entities. This could lead to increased social disparity and potentially exacerbate economic inequalities [56], particularly if low-income individuals are involved, similar to the situation observed in Axie Infinity [10].
The study of the emerging phenomena in novel platforms such as Decentraland and games such as Ice Poker is increasingly important. A better understanding of their underlying mechanisms and dynamics will allow stakeholders to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years. In particular, we invite policymakers to take a critical look into the growing phenomenon of gambling and gambling-like activities in social virtual worlds based on tokenomics. Given the risk-taking behavior associated with both cryptocurrencies and gambling, together with their novel use in immersive virtual worlds, consumer protection should be a priority, especially for vulnerable individuals, such as those who are both low-income and younger. To this end, policymakers could follow an approach similar to the recent tackling of abuse of loot boxes in video games [54], by expanding the legal definition of gambling and adapting existing regulation, such as enforcing the disclosure of probabilities regarding risks and rewards, and applying age rating and warning labels.
In view of the above legal and economic uncertainties, we also suggest developers of social virtual worlds to be wary of heavily relying —directly or indirectly— on such gambling-like activities to keep users engaged. Moreover, particular care should be taken in designing and implementing the decentralized governance and affordances of the platform, as to not allow a single third party to gain an overtly dominant role. Indeed, Decentraland’s reliance on ICE Poker represents a predicament for the platform, which already suffers from a volatile ecosystem and diminished interest, raising doubts on its long-term stability. In this regard, at the end of 2024 the project announced a new version of the platform, Decentraland 2.0, with the stated goals of strengthening social interactions, boosting user engagement, improving discoverability, and increased immersion and fun. At the moment of writing, it remains unclear if this new version will succeed in reaching its goals and if it will overcome the dominance of in-world gambling activities.
D.   Limitations